# What MLLMs Learn about When they Learn about Multimodal Reasoning: Perception, Reasoning, or their Integration?

### v2

@@ C.2 MathLens-General
Problems are sourced from six established multimodal reasoning datasets: BLINK (Fu et al. (2024)), V* (Wu & Xie (2024)), SpatialEval-Real (Wang et al. (2024a)), MMMU-Pro (Yue et al. (2024b)), EMMA (Hao et al. (2025)), and MathVista (Lu et al. (2024)), with mathematical geometry items excluded using metadata. We retain only multiple-choice problems to maintain consistency. Problems requiring more than two images are discarded, and dual-image inputs are concatenated into single images to ensure compatibility with models that do not support multi-image inputs.
@@
Problems that appear multimodal may in fact be solvable from text-only correlations (Yue et al. (2024b)), while others that look complex may reduce to simple pattern matching. To guard against such shortcuts, we use model-based validation to test multimodality and reasoning requirements. Each problem is evaluated with Gemini-2.5-Flash (Comanici et al. (2025)) under three settings: (1) full input, (2) text-only input without the image, and (3) full input with reasoning disabled. We generate eight responses per setting at temperature 0.6 to capture variability.
@@ Appendix D Implementation Details and resources
Details of the model configurations and corresponding sources are provided in Table 1.
@@
Model configuration for Figure 1.
@@
We use Eureka ML Insights Framework (Balachandran et al. (2024)) for reproducible evaluation. We run 7–9B parameter models on four NVIDIA A100 80GB GPUs, and 72B models on eight. Generation is accelerated and parallelized with the vLLM (Kwon et al. (2023)) library. Most experiments use greedy decoding (temperature $0.0$) for deterministic outputs. For models that otherwise suffer from text degeneration through severe repetition (e.g., Huang et al. (2025)), we apply stochastic decoding with temperature $0.6$ and top-$p$ $0.65$. The default maximum generation length is 32,768 tokens to accommodate long reasoning chains. For OVR (Wei et al. (2025)), we extend this limit to 48,000 tokens due to frequent truncations at lower cutoffs.
@@
Visual icons used in Footnote 2 are adapted from flaticon.com.
@@ E.1 Preliminaries
Prior work (Sun et al. (2025)) highlights that multimodal SFT datasets are considerably easier than textual SFT datasets, which contributes to weaker textual reasoning capacity in trained models. For example, the Vision-R1 dataset (Huang et al. (2025)) averages 821.5 tokens per reasoning trace with a 96.0% pass rate, whereas the text-only DeepMath dataset (He et al. (2025)) averages 8,207.8 tokens with a 75.0% pass rate. This difference suggests that multimodal SFT data require substantially less reasoning effort. Consistent with this, multimodal SFT models trained on such data underperform on standard reasoning benchmarks compared to textual SFT models. Finally, while textual SFT data can be constructed by distilling reasoning traces from large language models (Guo et al. (2025)), no comparably strong multimodal reasoning models with open reasoning traces currently exist, making effective multimodal SFT data generation particularly challenging.
@@ E.2 Further Insights
E.3 MathLens-General
Figure 14: Error type distribution across models on MathLens-General. Most shifts are associated with perception-related cases.
Figure 14 presents the error-type distribution for MathLens-General. Consistent with the main experiment (Section 3.4), most effects of multimodal reasoning finetuning concentrate on perception-related cases (Perception & Reasoning or Perception). However, these gains are less stable, reflecting that MathLens-General spans broader domains than MathLens and often demands out-of-distribution generalization from the multimodal training sets. An exception is textual SFT models, which show substantial reductions in pure Reasoning errors. This indicates that, unlike in math geometry tasks, the diverse reasoning skills required for MathLens-General are not well represented in the backbone (Qwen-2.5-VL). Finally, the higher fraction of Trivial correct cases arises from MathLens-General ’s multiple-choice format, in contrast to the open-ended geometry subset MathLens.
E.4 Full Results
@@
Table 2: Full results of all models used to produce the correlation plot in Figure 5 (left).
@@
Table 2 reports the full benchmark scores for all models used in the correlation analysis of Figure 5 (left). Results for MathVista, MathVerse, and MathVision are drawn from prior work (Wang et al. (2025)). Consequently, the set of models differs from our main evaluation and includes additional variants such as MM-Eureka (Meng et al. (2025)) and ThinkLite-VL (Deng et al. (2025)). These models were excluded from the main analysis for two reasons: (i) to maintain a balanced number of models across categories, particularly those trained with direct RL, and (ii) to focus on the stronger-performing models on other benchmarks.
Error type analysis.
Figure 15 show full error type analysis results for all models tested in this work.
@@
Table 3 reports the complete downstream evaluation results on MathLens, including all diagram modifications. These values underlie Figure 4Figure 5 (right),  Figure 8, and Figure 7. Note that the error-type analysis in Figure 15 relies on per-sample categorization and cannot be obtained directly from the aggregate scores presented here.
@@
Table 4 reports the full benchmark scores for each perception probe question type, providing the numerical values underlying Figure 9.
@@ Appendix F Qualitative Examples
We compare the fine-tuned multimodal reasoners with their corresponding backbone MLLMs. Figure 16 and Figure 17 present cases where fine-tuning corrected an initially wrong answer, while Figure 18 and Figure 19 illustrate cases where the model produced incorrect answers both before and after fine-tuning.
@@
Figure 20Figure 21, and Figure 22 show cases where textual SFT corrected initially wrong answers. Consistent with the quantitative results in Section 3.3, the stronger cognitive patterns induced by textual SFT also promote improved perception.

### v3

@@ C.2 MathLens-NS
Problems are sourced from six established multimodal reasoning datasets: BLINK ((Fu et al., 2024)), V* ((Wu & Xie, 2024)), SpatialEval-Real ((Wang et al., 2024a)), MMMU-Pro ((Yue et al., 2024b)), EMMA ((Hao et al., 2025)), and MathVista ((Lu et al., 2024)), with mathematical geometry items excluded using metadata. We retain only multiple-choice problems to maintain consistency. Problems requiring more than two images are discarded, and dual-image inputs are concatenated into single images to ensure compatibility with models that do not support multi-image inputs.
@@
Problems that appear multimodal may in fact be solvable from text-only correlations ((Yue et al., 2024b)), while others that look complex may reduce to simple pattern matching. To guard against such shortcuts, we use model-based validation to test multimodality and reasoning requirements. Each problem is evaluated with Gemini-2.5-Flash ((Comanici et al., 2025)) under three settings: (1) full input, (2) text-only input without the image, and (3) full input with reasoning disabled. We generate eight responses per setting at temperature 0.6 to capture variability.
@@ Appendix D Implementation Details and resources
Details of the model configurations and corresponding sources are provided in Table˜1.
@@
Model configuration for Figure˜1.
@@
We use Eureka ML Insights Framework ((Balachandran et al., 2024)) for reproducible evaluation. We run 7–9B parameter models on four NVIDIA A100 80GB GPUs, and 72B models on eight. Generation is accelerated and parallelized with the vLLM ((Kwon et al., 2023)) library. Most experiments use greedy decoding (temperature $0.0$) for deterministic outputs. For models that otherwise suffer from text degeneration through severe repetition (e.g., (Huang et al., 2025)), we apply stochastic decoding with temperature $0.6$ and top-$p$ $0.65$. The default maximum generation length is 32,768 tokens to accommodate long reasoning chains. For OVR ((Wei et al., 2025)), we extend this limit to 48,000 tokens due to frequent truncations at lower cutoffs.
@@
Visual icons used in Footnote˜2 are adapted from flaticon.com.
@@ E.1 Preliminaries
Prior work ((Sun et al., 2025)) highlights that multimodal SFT datasets are considerably easier than textual SFT datasets, which contributes to weaker textual reasoning capacity in trained models. For example, the Vision-R1 dataset ((Huang et al., 2025)) averages 821.5 tokens per reasoning trace with a 96.0% pass rate, whereas the text-only DeepMath dataset ((He et al., 2025)) averages 8,207.8 tokens with a 75.0% pass rate. This difference suggests that multimodal SFT data require substantially less reasoning effort. Consistent with this, multimodal SFT models trained on such data underperform on standard reasoning benchmarks compared to textual SFT models. Finally, while textual SFT data can be constructed by distilling reasoning traces from large language models ((Guo et al., 2025)), no comparably strong multimodal reasoning models with open reasoning traces currently exist, making effective multimodal SFT data generation particularly challenging.
@@ E.2 Further Insights
E.3 Full Results
Perception probe results by relation type.
We report perception probe performance decomposed by question type in Figure˜13. Each category groups probes according to the geometric relation being tested (e.g., angle relations, length relations, polygon structure). Performance gains differ substantially across categories. Angle-related relations and cocircularity show consistent improvement under multimodal reinforcement learning, while line-length relations exhibit limited change across training variants. Polygon-level relations, including triangle and quadrilateral detection, improve primarily when textual supervised fine-tuning is included. These differences appear correlated with the visual presentation of the target cue. Relations whose defining annotations are colocated with the relevant geometric element tend to benefit more from multimodal training, whereas relations relying on spatially separated marks or distant comparisons remain challenging.
@@ E.3 Full Results
Table 2: Full results of all models used to produce the correlation plot in Figure˜5.
@@
Table 2 reports the full benchmark scores for all models used in the correlation analysis of Figure˜5. Results for MathVista, MathVerse, and MathVision are drawn from prior work ((Wang et al., 2025)). Consequently, the set of models differs from our main evaluation and includes additional variants such as MM-Eureka ((Meng et al., 2025)) and ThinkLite-VL ((Deng et al., 2025)). These models were excluded from the main analysis for two reasons: (i) to maintain a balanced number of models across categories, particularly those trained with direct RL, and (ii) to focus on the stronger-performing models on other benchmarks.
@@
Table˜3 reports the complete downstream evaluation results on MathLens, including all diagram modifications. These values underlie Figure˜4Figure˜6,  Figure˜9, and Figure˜10. Note that the error-type analysis in Figure˜14 relies on per-sample categorization and cannot be obtained directly from the aggregate scores presented here.
@@
Table˜4 reports the full benchmark scores for each perception probe question type, providing the numerical values underlying Figure˜13.
Error type analysis.
Figure˜14 show full error type analysis results for all models tested in this work.
E.4 MathLens-NS
Figure 15: Data samples from MathLens-NS. We curate problem instance from a prior dataset ((Fu et al., 2024)) and annotate the text description and perception-centric question-answers.
Figure 16: Data samples from MathLens-NS. We curate problem instance from a prior dataset ((Hao et al., 2025)) and annotate the text description and perception-centric question-answers.
Figure 17: Error type distribution across models on MathLens-NS. Most shifts are associated with perception-related cases.
Figure˜17 presents the error-type distribution for MathLens-NS. Consistent with the main experiment (Section˜3.4), most effects of multimodal reasoning finetuning concentrate on perception-related cases (Perception & Reasoning or Perception). However, these gains are less stable, reflecting that MathLens-NS spans broader domains than MathLens and often demands out-of-distribution generalization from the multimodal training sets. An exception is textual SFT models, which show substantial reductions in pure Reasoning errors. This indicates that, unlike in math geometry tasks, the diverse reasoning skills required for MathLens-NS are not well represented in the backbone (Qwen-2.5-VL). Finally, the higher fraction of Trivial correct cases arises from MathLens-NS ’s multiple-choice format, in contrast to the open-ended geometry subset MathLens.
@@ Appendix F Qualitative Examples
We compare the fine-tuned multimodal reasoners with their corresponding backbone MLLMs. Figure˜18 and Figure˜19 present cases where fine-tuning corrected an initially wrong answer, while Figure˜20 and Figure˜21 illustrate cases where the model produced incorrect answers both before and after fine-tuning.
@@
Figure˜22Figure˜23, and Figure˜24 show cases where textual SFT corrected initially wrong answers. Consistent with the quantitative results in Section˜3.3, the stronger cognitive patterns induced by textual SFT also promote improved perception.